# Breaking the Dimensional Barrier: Dynamic Portfolio Choice with Parameter Uncertainty via Pontryagin Projection

### v2

@@ Appendix C Proof of Theorem 2
$p_{k}^{\Delta t,\theta}:=\mathbb{E}\!\left[p_{k}^{\mathrm{pw},\theta}\mid G_{t_{k}}^{\theta}\right],\hskip 18.49988ptp_{x,k}^{\Delta t,\theta}:=\mathbb{E}\!\left[p_{x,k}^{\mathrm{pw},\theta}\mid G_{t_{k}}^{\theta}\right],\hskip 18.49988ptp_{y,k}^{\Delta t,\theta}:=\mathbb{E}\!\left[p_{y,k}^{\mathrm{pw},\theta}\mid G_{t_{k}}^{\theta}\right].$
These are $G_{t_{k}}^{\theta}$-measurable and are the discrete-time counterparts of the $\theta$-conditional Pontryagin objects.
@@
(c) Discrete BSDE form and identification of the driver. Substituting the projection in (a) into the one-step BPTT recursion from Step 2 and then taking conditional expectations with respect to $G_{t_{k}}^{\theta}$ yields a canonical discrete-time BSDE representation for the adapted triple $(p_{k}^{\Delta t,\theta},z_{k}^{\theta},\tilde{z}_{k}^{\theta})$. The resulting drift coincides with the Euler discretization of the $\theta$-conditional adjoint BSDE associated with (45)–(46) under the fixed policy $\pi_{\varphi}$, up to standard one-step remainder terms of higher order in $\Delta t$. The same argument applies to the blocks $(p_{x,k}^{\Delta t,\theta},p_{y,k}^{\Delta t,\theta})$: they satisfy linearized discrete backward recursions (obtained by differentiating the discrete adjoint equations) and hence admit analogous discrete-BSDE representations with coefficients uniformly controlled on $K$. All estimates and remainder controls are identical to Huh et al. (2025b) once we condition on $\theta$, and the constants can be chosen uniformly on $K$ by the uniform-in-$\theta$ assumptions from Step 1.
Step 4: Passage to continuous time and convergence to the $\theta$-conditional PMP costate. For each fixed $\theta\in K$, the forward SDE and the $\theta$-conditional adjoint BSDE form a standard FBSDE with coefficients parametrized by $\theta$. Let $(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta})$ denote the continuous-time $\theta$-conditional Pontryagin objects under $\pi_{\varphi}$, with terminal condition $p_{T}^{\theta}=U^{\prime}(X_{T}^{\pi,\theta})$.
Define the piecewise-constant interpolations of the adapted discrete variables:
$p_{t}^{\Delta t,\theta}:=p_{k}^{\Delta t,\theta},\qquad p_{x,t}^{\Delta t,\theta}:=p_{x,k}^{\Delta t,\theta},\qquad p_{y,t}^{\Delta t,\theta}:=p_{y,k}^{\Delta t,\theta},\hskip 18.49988ptt\in[t_{k},t_{k+1}).$
@@
Then, by the same stability and convergence arguments as in Huh et al. (2025b) (Euler convergence for the forward equation plus convergence of the discrete-BSDE scheme for the backward equation), we obtain, for each fixed $\theta\in K$,
@@
as $\Delta t\to 0$. Because all Lipschitz, growth, ellipticity, and block-covariance constants were assumed uniform on $K$, the constants in these convergence bounds can be chosen independently of $\theta\in K$. This yields the claimed BPTT–PMP correspondence conditionally on $\theta$ and uniformly over $\theta$ in compact subsets of $\Theta$, completing the proof. ∎
Appendix D Auxiliary results for Theorem 3
D.1 Stability of the projection map $(A,G)\mapsto-A^{-1}G$
@@
Let $D$ be a measurable domain and let $\mu$ be a reference measure on $D$. Let $A,\widetilde{A}:D\to\mathbb{R}^{d\times d}$ and $G,\widetilde{G}:D\to\mathbb{R}^{d}$ be measurable. Assume:
@@
    $A(z)$ is invertible for $\mu$-a.e. $z\in D$ and $\|A^{-1}\|_{L^{\infty}(D)}\leq\kappa$ for some $\kappa>0$;
@@
    $\|\widetilde{A}-A\|_{L^{\infty}(D)}\leq(2\kappa)^{-1}$.
Define $\pi:=-A^{-1}G$ and $\widetilde{\pi}:=-\widetilde{A}^{-1}\widetilde{G}$. Then $\widetilde{A}(z)$ is invertible for $\mu$-a.e. $z\in D$ with $\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\leq 2\kappa$, and
$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\;\leq\;2\kappa\,\|\widetilde{G}-G\|_{L^{2}(\mu)}\;+\;2\kappa^{2}\bigl(M+\|\widetilde{G}\|_{L^{\infty}(D)}\bigr)\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$ (95)
@@
Throughout, $\|\cdot\|$ denotes the operator norm induced by the Euclidean norm.
@@
$\widetilde{A}(z)\;=\;A(z)\Bigl(I+A(z)^{-1}\bigl(\widetilde{A}(z)-A(z)\bigr)\Bigr).$
By (i) and (iii),
$\bigl\|A^{-1}(z)\bigl(\widetilde{A}(z)-A(z)\bigr)\bigr\|\;\leq\;\|A^{-1}\|_{L^{\infty}(D)}\,\|\widetilde{A}-A\|_{L^{\infty}(D)}\;\leq\;\kappa\cdot(2\kappa)^{-1}\;=\;\tfrac{1}{2}.$
Hence $I+A^{-1}(\widetilde{A}-A)$ is invertible and $\bigl\|\bigl(I+A^{-1}(\widetilde{A}-A)\bigr)^{-1}\bigr\|\leq(1-\tfrac{1}{2})^{-1}=2$. Therefore $\widetilde{A}(z)$ is invertible and
$\|\widetilde{A}^{-1}(z)\|\;=\;\bigl\|\bigl(I+A^{-1}(\widetilde{A}-A)\bigr)^{-1}A^{-1}(z)\bigr\|\;\leq\;2\|A^{-1}\|_{L^{\infty}(D)}\;\leq\;2\kappa,$
so $\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\leq 2\kappa$.
@@
$\|\widetilde{A}^{-1}-A^{-1}\|_{L^{2}(\mu)}\;\leq\;\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\,\|\widetilde{A}-A\|_{L^{2}(\mu)}\,\|A^{-1}\|_{L^{\infty}(D)}\;\leq\;2\kappa^{2}\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$
@@
$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\leq\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\,\|\widetilde{G}-G\|_{L^{2}(\mu)}+\|\widetilde{A}^{-1}-A^{-1}\|_{L^{2}(\mu)}\,\|G\|_{L^{\infty}(D)}.$
@@
$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\leq 2\kappa\,\|\widetilde{G}-G\|_{L^{2}(\mu)}+2\kappa^{2}\,M\,\|\widetilde{A}-A\|_{L^{2}(\mu)}\leq 2\kappa\,\|\widetilde{G}-G\|_{L^{2}(\mu)}+2\kappa^{2}\bigl(M+\|\widetilde{G}\|_{L^{\infty}(D)}\bigr)\|\widetilde{A}-A\|_{L^{2}(\mu)},$
@@
D.2 Slab-wise small-gain for the $q$-aggregated projection inputs
@@
Assume the working domain carries a time coordinate and, for concreteness, $D\subset[0,T]\times S$ and $\mu(dt,d\xi)=dt\otimes\nu(d\xi)$ for some reference measure $\nu$ on $S$. Fix a partition $0=t_{0}<t_{1}<\cdots<t_{K}=T$ with slab lengths $\tau_{k}:=t_{k}-t_{k-1}$ and define
$D_{k}:=D\cap([t_{k-1},t_{k}]\times S),\hskip 18.49988pt\mu_{k}:=\mu|_{D_{k}},\hskip 18.49988pt\|f\|_{k}:=\|f\|_{L^{2}(\mu_{k})}.$
Then $\|f\|_{L^{2}(\mu)}^{2}=\sum_{k=1}^{K}\|f\|_{k}^{2}$.
@@
Let $U$ be a neighborhood of a locally optimal deployable $\theta$-blind policy $\pi^{\star}_{\mathrm{blind}}$ in $L^{2}(\mu)$ such that for all $\pi\in U$,
$\|A_{\pi}^{-1}\|_{L^{\infty}(D)}\leq\kappa,\hskip 18.49988pt\|G^{\mathrm{mix}}_{\pi}\|_{L^{\infty}(D)}\leq M_{G}$
for some constants $\kappa,M_{G}>0$. Assume that on each slab $D_{k}$ the $q$-aggregated projection inputs satisfy
$\|A_{\pi_{1}}-A_{\pi_{2}}\|_{k}\leq\overline{L}_{A}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},\hskip 18.49988pt\|G^{\mathrm{mix}}_{\pi_{1}}-G^{\mathrm{mix}}_{\pi_{2}}\|_{k}\leq\overline{L}_{G}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},$ (96)
for all $\pi_{1},\pi_{2}\in U$, with constants $\overline{L}_{A},\overline{L}_{G}>0$ depending only on band data. Define the (population) projection map $T(\pi):=-A_{\pi}^{-1}G^{\mathrm{mix}}_{\pi}$ and
$\rho(\tau):=\Bigl(\kappa\,\overline{L}_{G}+\kappa^{2}M_{G}\,\overline{L}_{A}\Bigr)\tau^{1/2}.$
Then for each slab $D_{k}$ and all $\pi_{1},\pi_{2}\in U$,
$\|T(\pi_{1})-T(\pi_{2})\|_{k}\leq\rho(\tau_{k})\,\|\pi_{1}-\pi_{2}\|_{k}.$
@@
$\rho^{\ast}:=\max_{1\leq k\leq K}\rho(\tau_{k})<1,$
then $T$ is a strict contraction on every slab with constant at most $\rho^{\ast}$.
@@
Fix $k$ and $\pi_{1},\pi_{2}\in U$. Set $A_{i}:=A_{\pi_{i}}$ and $G_{i}:=G^{\mathrm{mix}}_{\pi_{i}}$. Then
$T(\pi_{1})-T(\pi_{2})=-A_{1}^{-1}(G_{1}-G_{2})-(A_{1}^{-1}-A_{2}^{-1})G_{2},\hskip 18.49988ptA_{1}^{-1}-A_{2}^{-1}=A_{1}^{-1}(A_{2}-A_{1})A_{2}^{-1}.$
Using Hölder ($L^{\infty}\times L^{2}\to L^{2}$) on $D_{k}$ together with $\|A_{i}^{-1}\|_{L^{\infty}(D)}\leq\kappa$ and $\|G_{2}\|_{L^{\infty}(D)}\leq M_{G}$ yields
$\|T(\pi_{1})-T(\pi_{2})\|_{k}\leq\kappa\|G_{1}-G_{2}\|_{k}+\kappa^{2}M_{G}\|A_{1}-A_{2}\|_{k}.$
Applying (96) gives the claim with $\rho(\tau_{k})$. ∎
Remark 3 (Where the $\tau^{1/2}$ gain comes from; relation to huh2025a).
@@
A full PDE derivation of this short-time contraction philosophy is given in Huh et al. (2025b, Appendix B); see also the companion supplementary note to Huh et al. (2025b) (“On Proof of Theorem 4 (Policy Gap Bound)”) for a step-by-step slab contraction argument in a closely related setting. In the present paper the same smoothing principle is applied not to the costate update map directly, but to the induced $q$-aggregated projection inputs $(A_{\pi},G^{\mathrm{mix}}_{\pi})$.
D.3 Proof of Theorem 3

### v3

@@ Appendix D Proof of Theorem 2
$p_{k}^{\Delta t,\theta}:=\mathbb{E}\!\left[p_{k}^{\mathrm{pw},\theta}\mid\mathcal{G}_{t_{k}}^{\theta}\right],\qquad p_{x,k}^{\Delta t,\theta}:=\mathbb{E}\!\left[p_{x,k}^{\mathrm{pw},\theta}\mid\mathcal{G}_{t_{k}}^{\theta}\right],\qquad p_{y,k}^{\Delta t,\theta}:=\mathbb{E}\!\left[p_{y,k}^{\mathrm{pw},\theta}\mid\mathcal{G}_{t_{k}}^{\theta}\right].$
@@
By construction, $(p_{k}^{\Delta t,\theta},z_{k}^{\theta},\tilde{z}_{k}^{\theta})$ satisfy a discrete-time BSDE representation. Moreover, substituting the projection into the one-step BPTT recursion and taking conditional expectations yields that the discrete driver coincides with the Euler discretization of the $\theta$-conditional adjoint BSDE under the fixed policy $\pi_{\varphi}$, up to standard one-step remainder terms controlled by Taylor/Euler estimates. Under the standing smoothness assumptions, the projection residual $\varepsilon_{k+1}^{\theta}$ contributes only higher-order discretization error; all bounds can be chosen uniformly for $\theta\in\Theta_{0}$ (as in Huh et al. (2025)).
Step 4: Convergence as $\Delta t\to 0$ and uniformity on compacts. Let $(p_{t}^{\theta},z_{t}^{\theta},\tilde{z}_{t}^{\theta})$ denote the continuous-time $\theta$-conditional adjoint triple under $\pi_{\varphi}$, and let $(p_{x,t}^{\theta},p_{y,t}^{\theta})$ denote the corresponding sensitivity blocks (in smooth Markov regimes, value-derivative blocks).
Define the piecewise-constant interpolations of the adapted discrete variables on $[0,T]$:
$p_{t}^{\Delta t,\theta}:=p_{k}^{\Delta t,\theta},\qquad p_{x,t}^{\Delta t,\theta}:=p_{x,k}^{\Delta t,\theta},\qquad p_{y,t}^{\Delta t,\theta}:=p_{y,k}^{\Delta t,\theta},\qquad z_{t}^{\Delta t,\theta}:=z_{k}^{\theta},\qquad\tilde{z}_{t}^{\Delta t,\theta}:=\tilde{z}_{k}^{\theta},\hskip 18.49988ptt\in[t_{k},t_{k+1}).$
@@
Standard stability of Euler for the forward SDE together with convergence of the corresponding discrete BSDE scheme implies, for each fixed $\theta\in\Theta_{0}$,
@@
and, for the martingale coefficients,
$\mathbb{E}\!\left[\int_{0}^{T}\|z_{t}^{\Delta t,\theta}-z_{t}^{\theta}\|^{2}\,dt\right]\to 0,\hskip 18.49988pt\mathbb{E}\!\left[\int_{0}^{T}\|\tilde{z}_{t}^{\Delta t,\theta}-\tilde{z}_{t}^{\theta}\|^{2}\,dt\right]\to 0,$
as $\Delta t\to 0$. Because all Lipschitz, growth, nondegeneracy, and smoothness constants were assumed uniform on $\Theta_{0}$, the constants in these convergence bounds can be chosen independently of $\theta\in\Theta_{0}$. This yields the claimed BPTT–PMP correspondence conditionally on $\theta$ and uniformly over $\theta$ in compact subsets of $\Theta$. ∎
Appendix E Auxiliary results for Theorem 3
E.1 Stability of the projection map $(A,g)\mapsto-A^{-1}g$
@@
Let $\mathcal{D}$ be a measurable domain and let $\mu$ be a reference measure on $\mathcal{D}$. Let $A,\widetilde{A}:\mathcal{D}\to\mathbb{R}^{d\times d}$ and $g,\widetilde{g}:\mathcal{D}\to\mathbb{R}^{d}$ be measurable. Assume:
@@
    $A(z)$ is invertible for $\mu$-a.e. $z\in\mathcal{D}$ and $\|A^{-1}\|_{L^{\infty}(\mathcal{D})}\leq\kappa$ for some $\kappa>0$;
@@
    $\|\widetilde{A}-A\|_{L^{\infty}(\mathcal{D})}\leq(2\kappa)^{-1}$.
Define $\pi:=-A^{-1}g$ and $\widetilde{\pi}:=-\widetilde{A}^{-1}\widetilde{g}$. Then $\widetilde{A}(z)$ is invertible for $\mu$-a.e. $z\in\mathcal{D}$ with $\|\widetilde{A}^{-1}\|_{L^{\infty}(\mathcal{D})}\leq 2\kappa$, and
$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\;\leq\;2\kappa\,\|\widetilde{g}-g\|_{L^{2}(\mu)}\;+\;2\kappa^{2}M_{g}\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$ (100)
@@
Throughout, $\|\cdot\|$ denotes the operator norm induced by the Euclidean norm, and $\|\cdot\|_{L^{\infty}(\mathcal{D})}$ denotes the $\mu$-essential supremum of the corresponding pointwise norm on $\mathcal{D}$.
@@
$\widetilde{A}(z)\;=\;A(z)\Bigl(I_{d}+A(z)^{-1}\bigl(\widetilde{A}(z)-A(z)\bigr)\Bigr),$
where $I_{d}$ is the $d\times d$ identity matrix. By (i) and (iii),
$\bigl\|A^{-1}(z)\bigl(\widetilde{A}(z)-A(z)\bigr)\bigr\|\;\leq\;\|A^{-1}\|_{L^{\infty}(\mathcal{D})}\,\|\widetilde{A}-A\|_{L^{\infty}(\mathcal{D})}\;\leq\;\kappa\cdot(2\kappa)^{-1}\;=\;\tfrac{1}{2}.$
Hence $I_{d}+A^{-1}(\widetilde{A}-A)$ is invertible and $\bigl\|\bigl(I_{d}+A^{-1}(\widetilde{A}-A)\bigr)^{-1}\bigr\|\leq(1-\tfrac{1}{2})^{-1}=2$. Therefore $\widetilde{A}(z)$ is invertible and
$\|\widetilde{A}^{-1}(z)\|\;=\;\bigl\|\bigl(I_{d}+A^{-1}(\widetilde{A}-A)\bigr)^{-1}A^{-1}(z)\bigr\|\;\leq\;2\|A^{-1}\|_{L^{\infty}(\mathcal{D})}\;\leq\;2\kappa,$
so $\|\widetilde{A}^{-1}\|_{L^{\infty}(\mathcal{D})}\leq 2\kappa$.
@@
$\|\widetilde{A}^{-1}-A^{-1}\|_{L^{2}(\mu)}\;\leq\;\|\widetilde{A}^{-1}\|_{L^{\infty}(\mathcal{D})}\,\|\widetilde{A}-A\|_{L^{2}(\mu)}\,\|A^{-1}\|_{L^{\infty}(\mathcal{D})}\;\leq\;2\kappa^{2}\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$
@@
$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\leq\|\widetilde{A}^{-1}\|_{L^{\infty}(\mathcal{D})}\,\|\widetilde{g}-g\|_{L^{2}(\mu)}+\|\widetilde{A}^{-1}-A^{-1}\|_{L^{2}(\mu)}\,\|g\|_{L^{\infty}(\mathcal{D})}.$
@@
$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\leq 2\kappa\,\|\widetilde{g}-g\|_{L^{2}(\mu)}+2\kappa^{2}\,M_{g}\,\|\widetilde{A}-A\|_{L^{2}(\mu)},$
@@
E.2 Slab-wise small-gain for the $q$-aggregated projection inputs
@@
Assume the working domain carries a time coordinate and, for concreteness, $\mathcal{D}\subset[0,T]\times\mathcal{S}$ and $\mu(dt,d\xi)=dt\otimes\nu(d\xi)$ for some reference measure $\nu$ on $\mathcal{S}$. Fix a partition $0=t_{0}<t_{1}<\cdots<t_{L}=T$ with slab lengths $\tau_{k}:=t_{k}-t_{k-1}$ and define
$\mathcal{D}_{k}:=\mathcal{D}\cap([t_{k-1},t_{k}]\times\mathcal{S}),\hskip 18.49988pt\mu_{k}:=\mu|_{\mathcal{D}_{k}},\hskip 18.49988pt\|f\|_{k}:=\|f\|_{L^{2}(\mu_{k})}.$
Then $\|f\|_{L^{2}(\mu)}^{2}=\sum_{k=1}^{L}\|f\|_{k}^{2}$.
@@
Let $\mathcal{U}$ be a subset of the admissible policy class that is an $L^{2}(\mu)$-neighborhood of a locally optimal deployable $\theta$-blind policy $\pi^{\star,\mathrm{blind}}$ (in particular $\pi^{\star,\mathrm{blind}}\in\mathcal{U}$), and on which the following uniform bounds hold for all $\pi\in\mathcal{U}$:
$\|A_{\pi}^{-1}\|_{L^{\infty}(\mathcal{D})}\leq\kappa,\hskip 18.49988pt\|g^{\mathrm{mix}}_{\pi}\|_{L^{\infty}(\mathcal{D})}\leq M_{g}$
for some constants $\kappa,M_{g}>0$. Assume that on each slab $\mathcal{D}_{k}$ the $q$-aggregated projection inputs satisfy
$\|A_{\pi_{1}}-A_{\pi_{2}}\|_{k}\leq\bar{L}_{A}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},\hskip 18.49988pt\|g^{\mathrm{mix}}_{\pi_{1}}-g^{\mathrm{mix}}_{\pi_{2}}\|_{k}\leq\bar{L}_{g}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},$ (101)
for all $\pi_{1},\pi_{2}\in\mathcal{U}$, with constants $\bar{L}_{A},\bar{L}_{g}>0$ depending only on band data. Define the (population) projection map $\mathcal{T}(\pi):=-A_{\pi}^{-1}g^{\mathrm{mix}}_{\pi}$ and
$\varrho(\tau):=\Bigl(\kappa\,\bar{L}_{g}+\kappa^{2}M_{g}\,\bar{L}_{A}\Bigr)\tau^{1/2}.$
Then for each slab $\mathcal{D}_{k}$ and all $\pi_{1},\pi_{2}\in\mathcal{U}$,
$\|\mathcal{T}(\pi_{1})-\mathcal{T}(\pi_{2})\|_{k}\leq\varrho(\tau_{k})\,\|\pi_{1}-\pi_{2}\|_{k}.$
@@
$\varrho_{*}:=\max_{1\leq k\leq L}\varrho(\tau_{k})<1,$
then $\mathcal{T}$ is a strict contraction on every slab with constant at most $\varrho_{*}$.
@@
Fix $k$ and $\pi_{1},\pi_{2}\in\mathcal{U}$. Set $A_{i}:=A_{\pi_{i}}$ and $g_{i}:=g^{\mathrm{mix}}_{\pi_{i}}$. Then
$\mathcal{T}(\pi_{1})-\mathcal{T}(\pi_{2})=-A_{1}^{-1}(g_{1}-g_{2})-(A_{1}^{-1}-A_{2}^{-1})g_{2},\hskip 18.49988ptA_{1}^{-1}-A_{2}^{-1}=A_{1}^{-1}(A_{2}-A_{1})A_{2}^{-1}.$
Using Hölder ($L^{\infty}\times L^{2}\to L^{2}$) on $\mathcal{D}_{k}$ together with $\|A_{i}^{-1}\|_{L^{\infty}(\mathcal{D})}\leq\kappa$ and $\|g_{2}\|_{L^{\infty}(\mathcal{D})}\leq M_{g}$ yields
$\|\mathcal{T}(\pi_{1})-\mathcal{T}(\pi_{2})\|_{k}\leq\kappa\|g_{1}-g_{2}\|_{k}+\kappa^{2}M_{g}\|A_{1}-A_{2}\|_{k}.$
Applying (101) gives the claim with $\varrho(\tau_{k})$. ∎
Remark 6 (Where the $\tau^{1/2}$ gain comes from; relation to Huh et al. (2025)).
@@
A full PDE derivation of this short-time contraction philosophy is given in Huh et al. (2025, Appendix B); see also the companion supplementary note to Huh et al. (2025) (“On Proof of Theorem 4 (Policy Gap Bound)”) for a step-by-step slab contraction argument in a closely related setting. In the present paper the same smoothing principle is applied not to the costate update map directly, but to the induced $q$-aggregated projection inputs $(A_{\pi},g^{\mathrm{mix}}_{\pi})$.
E.3 Proof of Theorem 3